# Quantum Wave Simulation with Sources and Loss Functions

### v1

@@ 4.1 Single Pulse Source
Consider a single point-source defined by $\hat{\mathbf{s}}(\mathbf{x},t)=\hat{\boldsymbol{\chi}}\delta(\mathbf{x}-\mathbf{x}_{s})f(t)$. Its implementation complexity is tied to the number of values that need to be initialized on the QC. We identify two optimal scenarios:
4.1.1 Fixed Source Duration
@@
Since the required initialization volume does not scale with $N$, only a constant number of grid points needs to be initialized, ensuring $\mathcal{O}(1)$ complexity.
4.1.2 Increasing Grid Resolution
@@
4.2 Multiple Pulse Sources
We extend our analysis to sources composed of multiple contributions as defined in (27).
4.2.1 Synchronized Activation
@@ 4.2.2 Asynchronous Activation
$\phi(0)=(I^{\otimes\log_{2}S)}\otimes\mathbf{B}^{1/2})[\mathbf{w}_{1};\dots;\mathbf{w}_{S}],$ (29)
@@
This effectively time-dilates each component, aligning them to $T^{\text{sync}}$. Subsequently, the synchronized state is evolved under the multi-state wave equation Hamiltonian $\mathbf{H}^{\text{mult}}=I^{\otimes\log_{2}S}\otimes\mathbf{H}$ to the final simulation time $T$:
@@
Due to the linearity of the wave equation, the superposition of partial wave fields in |$\ket{\psi(T)}$ corresponds to the solution of (21) at time $T$. Consequently, we can estimate an $l_{2}$-distance to a target field or subspace energies using Theorem 2. Since the addition operation is inherently non-unitary, it must be integrated within the measurement process. Furthermore, as the loss function is nonlinear, we cannot compute losses for individual sub-fields separately and then combine them classically to derive the joint loss; therefore, we have to rely on Theorem 2 to estimate the loss function.
@@
4.3 General Sources
Implementing an arbitrary source time function $f(t)$ that lacks compact support in time is inefficient with the above approach, as it results in initial conditions lacking compact support in space, which are prohibitively expensive to compute classically and initialize on the QC. However, under certain regularity assumptions about the medium surrounding the source, we can exploit rotational symmetries in the wave fields during initialization. Specifically, initial conditions, which satisfy rotational covariance can be implemented efficiently:
@@ 4.3 General Sources
Let $\mathbf{K}:\mathbb{R}^{D}\to\mathbb{R}^{C}$, where $C\leq D$. Assume that $\mathbf{K}$ is rotationally covariant around a point $\mathbf{x}_{0}\in\mathbb{R}^{D}$ in the first $C$ dimensions and invariant in the remaining $D-C$ dimensions. Consider an isotropic radially symmetric grid centered at $\mathbf{x}_{0}$, discretized into $N=A^{D}$ points, where $A$ is the number of radial divisions. The grid points are specified by radial indices $a\in\{1,\dots,A\}$ and angular indices $\boldsymbol{\theta}$. Define the quantum state
|$\ket{\psi}=\frac{1}{\sqrt{\mathcal{N}}}\sum_{c=1}^{C}\sum_{a=1}^{A}\sum_{\boldsymbol{\theta}}K_{c}\big{(}\mathbf{x}_{a,\boldsymbol{\theta}}\big{)}\ket{c}\otimes\ket{a}\otimes\ket{\boldsymbol{\theta}},$ (35)
@@
where $K_{c}\big{(}\mathbf{x}_{a,\boldsymbol{\theta}}\big{)}$ is the $c^{\text{th}}$ component of $\mathbf{K}$ at grid point $\mathbf{x}_{a,\boldsymbol{\theta}}$, and $\mathcal{N}=\sum_{c,a,\boldsymbol{\theta}}\left|K_{c}\big{(}\mathbf{x}_{a,\boldsymbol{\theta}}\big{)}\right|^{2}$ ensures normalization. This state can be initialized with $\mathcal{O}(N^{1/D})$ classical evaluations of $\mathbf{K}$ and $\tilde{\mathcal{O}}(CN^{1/D})$ quantum gates, where the tilde hides polylogarithmic factors.
Examples for $\mathbf{K}$ include the $D$-dimensional acoustic pressure field ($C=1$), the acoustic particle velocity field, the electric field, and the magnetic field (all $C=D$). The elastic case is not included, as there $C>D$.
Assuming a non-vanishing volume of homogeneous and isotropic material up to radius $r_{s}$ around the point-source location we can use analytic Greens functions to map the source to symmetric initial conditions:
Theorem 5 (Source initialization, Appendix B).
Let $f:[0,T]\to\mathbb{R}$ be a compactly supported function, and let $\mathbf{G}:\mathbb{R}^{D}\times[0,\infty)\to\mathbb{R}^{C}$ be a rotationally covariant Green’s function around $\mathbf{x}_{0}$ in the first $C$ dimensions and invariant in the remaining dimensions; at least up to radius $r_{s}$, which is constant and non-vanishing. Define the time convolution $\mathbf{K}=f*\mathbf{G}$. Using an isotropic radially symmetric grid centered at $\mathbf{x}_{0}$ with $N=\mathcal{O}(A^{D})$ points, the discretized quantum state encoding of $\mathbf{K}$ can be initialized with $\mathcal{O}(N^{1/D})$ classical evaluations and $\tilde{\mathcal{O}}(CN^{1/D})$ quantum gates, where the tilde hides polylogarithmic factors.
As the homogeneous and isotropic volume around the source is assumed to be small this method employs windowing as detailed in Appendix B and necessitates a combination of wave fields in the read out process as discussed in Section 4.2.2. The presented approach also allows for contributions $\boldsymbol{\chi}_{s}(\mathbf{x})$ that are not point-sources, as long as their combined initial fields satisfy rotational covariance and if $\boldsymbol{\chi}_{s}(\mathbf{x})$ is contained in the homogeneous and isotropic volume. If $\boldsymbol{\chi}_{s}(\mathbf{x})$ is non-symmetric, it can in some cases be decomposed into symmetric parts, provided the number of parts does not scale with $N$, thus preserving the algorithm’s overall scaling. For multiple general source contributions (27), a separate spherical grid is required for each source. Bounding the error introduced in the loss function estimation by this approach is left for future research.
5 Runtime Advantage
Given that the discretized Hamiltonian $\mathbf{H}$ (8) is $d$-sparse, i.e. it has at most $d=\mathcal{O}(1)$ nonzero values in each row and column, which is ensured when $\mathbf{A}$ and $\mathbf{B}$ are $d$-sparse matrices, quantum simulation requires $\mathcal{O}(\log N)$ operations (Theorem 1). In principle this offers an exponential speed-up over classical solvers, which all scale at least with $\mathcal{O}(N)$ [49, 50, 51, 52, 53]. However, as we now show, the obtainable speed-up is usually only polynomial [35]. In wave-based problems, increasing grid points typically aims to (i) enlarge the physical domain or (ii) increase the simulated frequencies. For systems with local couplings, such as those resulting from discretizing the considered wave equations, enlarging the domain requires a longer simulation time $t$ for waves to reach the boundaries. Similarly, a finer discretization to simulate higher frequencies raises $\|\mathbf{H}\|_{\max}$ due to the reduced grid spacing. As per Theorem 1, the number of query calls to $\mathcal{H}$ scales linearly with both $t$ and $\|\mathbf{H}\|_{\max}$. Consequently, the quantum speed-up over classical algorithms becomes a polynomial factor dependent on the spatial dimension $D$. Specifically, let $t=\mathcal{O}(N^{1/D})$ represent the simulation time for the signal to reach the domain boundary in $D$ dimensions. For classical algorithms scaling linearly with $N$, the overall runtime complexity is $\mathcal{O}(Nt)=\mathcal{O}(N^{1+1/D})$. In contrast, the HS algorithm has runtime $\tilde{\mathcal{O}}(t)=\tilde{\mathcal{O}}(N^{1/D})$, where the tilde hides logarithmic factors. Thus, we expect a quadratic speed-up in 1D, cubic in 2D, and quartic in 3D. In contrast, [35] demonstrate an exponential speed-up for oscillator models with purely non-local couplings, where the simulation time does not scale with the system size.
Additionally, the algorithm’s oracle $\mathcal{H}$ imposes constraints: to preserve quantum speed-up, the information transferred to the QC must be of order $\tilde{\mathcal{O}}(N^{1/D})$. This limits the number of independent material parameters defining the propagation this order, necessitating functional or low-dimensional material representations.
6 Limitations
@@
6.1 Loss function Globality in Space and Compactness in Time
We identify two limitations within the introduced loss function estimations:
(i) Estimating time-dependent loss functions (as described in, e.g., [54]): Currently, our algorithm estimates the loss at a single measurement time step. Computing time-dependent loss functions involves separate simulation experiments for each measurement time, with the overall loss metric computed across all steps using either Theorem 2 or classical post-processing. However, this approach becomes inefficient for long comparison/integration times, where the number of comparison time steps scales worse than $\mathcal{O}(\text{polylog}(N))$. Non-destructive estimation of the loss function [55, 56] could provide a solution by preserving the state throughout the measurement operation, enabling iterative evolution and loss estimation of the wave field for multiple time points.
(ii) Globality of measurement quantities: A limitation of QC loss estimation via observable measurements is the requirement for sufficiently global measurement quantities. Suppose a state is measured in a subspace with only a small probability amplitude fraction compared to the entire quantum state. In that case, more samples are needed to estimate the norm accurately, as the expected value becomes vanishingly small. Ensuring the subspace size is large with respect to $N$ mitigates this issue.
6.2 Compactness of Source Terms
@@
(ii) Situations in which the number of initial conditions grows polynomially with $N$ must be avoided to prevent restricting the overall runtime. An example of this could be a continuously active source within a domain of increasing size and simulation time.
(iii) For a general source term a radial symmetric grid is required. This may be practically limiting as domains are often non-symmetric around the source location. Additionally, the introduced vector field initialization algorithm does not cover cases where the vector field dimension is larger than the spatial one.
6.3 Damping
A significant limitation of the presented framework is its inability to implement damping due to its non-unitary nature. However, damping in the context of HS has been recently studied, demonstrating that quantum speed-up can be preserved under certain conditions [57]. An alternative approach is to map a damped classical wave equation to an open quantum system. Simulating such systems on QCs is a recent research focus [58]. We believe that our work can eventually be combined with these approaches.
6.4 Time-Dependent Media
A final remark concerns time-dependent media, such as those in time-modulated meta-materials [e.g., 59, 60, 61]. In these cases, $\mathbf{B}=\mathbf{B}(t)$ becomes time-dependent, implying that it no longer commutes with the time derivative. Consequently, a time-dependent symmetric problem is mapped to a time-dependent non-Hermitian Schrödinger equation, rendering the introduced framework inapplicable [62].
7 Code
We present an educational Git repository, where we implement all the presented methods and concepts for the acoustic wave equation and Maxwell’s equations. The numerical implementation is based on a staggered grid finite difference method as detailed in Appendix C. The repository can be found here.
8 Conclusions
@@
Our method is compatible with any numerical approach that respects the anti-Hermitian nature of the operator and results in a diagonal or block diagonal representation of the material properties. The introduced representation of the wave equations naturally leads to a quantum encoding that enables the measurement of energies in subspaces of the model with optimal precision scaling. Additionally, we have shown that the $l_{2}$-distance of the final simulated wave field to a target or observed wave field can be measured with optimal precision scaling and that a broad class of source terms can be implemented efficiently.
Recent studies have indicated that HS could be feasible on noisy QCs, potentially preceding the advent of noise-free quantum hardware [63, 64]. Furthermore, new advances in quantum error correction hint that noise-free QCs might be realized soon [65]. These findings suggest that our proposed framework may have practical applications in the near future, thereby underscoring the importance of continued research in this field. Our future research will focus on oracle free and optimal encodings of the evolution Hamiltonian and the material transforms.
@@ Acknowledgements.
We would like to thank Dr. Scott Keating for helpful discussions on the loss functions and Marion Dugué for insights on the source decompositions. Furthermore, we are thankful to Patrick Marty and Dr. Ines Ulrich for sharing insights on acoustic wave physics in the context of medical imaging and to Dr. Václav Hapla for initial brainstorming sessions.
Grant Acknowledgment: C.B. was supported by the Swiss National Science Foundation through a Postdoc.Mobility Fellowship. Furthermore, we gratefully acknowledge the support provided by Google Quantum AI, which has significantly contributed to the advancement of this research.
@@
Code availability statement: All codes used in this work are available at
Author contribution statement: study conception: C. Bösch; sources: M. Schade and C. Bösch; loss functions and energy estimation: M. Schade and C. Bösch; acoustic FD implementation: G. Aloisi; general boundary conditions and constraints: G. Aloisi and C. Bösch; manuscript preparation: M. Schade, C. Bösch, G. Aloisi. All authors reviewed the results and approved the final version of the manuscript.

### v3

@@ V.1 Single Pulse Source
where $\hat{\pmb{\chi}}\in\mathbb{R}^{C}$ is an arbitrary vector. Its implementation complexity is tied to the number of values that need to be initialized on the QC. We identify two optimal scenarios:
V.1.1 Fixed Source Duration
@@
Since the required initialization volume does not scale with $N$, only a constant number of grid points needs to be initialized, ensuring $\mathcal{O}(1)$ runtime complexity.
V.1.2 Increasing Grid Resolution
@@
V.2 Multiple Pulse Sources
We extend our analysis to sources composed of multiple point sources
$\hat{\mathbf{s}}(\mathbf{x},t)=\sum_{s=1}^{S}\hat{\pmb{\chi}}_{s}\delta(\mathbf{x}-\mathbf{x}_{s})f_{s}(t).$ (28)
V.2.1 Synchronous Activation
@@ V.2.2 Asynchronous Activation
$\phi(0)=(I^{\otimes\log_{2}(S)}\otimes\mathbf{B}^{1/2})[\mathbf{w}_{1};\dots;\mathbf{w}_{S}],$ (29)
@@
This effectively time-dilates each component, aligning them to $T^{\text{sync}}$. Subsequently, the synchronized state is evolved under the multi-state wave equation Hamiltonian $\mathbf{H}^{\text{mult}}=I^{\otimes\log_{2}(S)}\otimes\mathbf{H}$ to the final simulation time $T$:
@@
Due to the linearity of the wave equation, the superposition of partial wave fields in |$\ket{\psi(T)}$ corresponds to the solution of Eq. (21) at time $T$. Consequently, we can estimate an $l_{2}$-distance to a target field or subspace energies using Theorem 2. Since the addition operation is inherently non-unitary, it must be integrated within the measurement process. Furthermore, as the loss function is nonlinear, we cannot compute losses for individual sub-fields separately and then combine them classically to derive the joint loss; therefore, we have to rely on Theorem 2 to estimate the loss function.
@@
V.3 Rotationally symmetric sources without compact support in time
Implementing a source time function $f(t)$ that lacks compact support in time is inefficient with the above approach, as it results in initial conditions lacking compact support in space. Such initial fields are prohibitively expensive to compute classically and initialize on the QC.
To target this, we introduce an algorithm that efficiently initializes rotationally covariant wave fields as quantum states:
@@ V.3 Rotationally symmetric sources without compact support in time
Let $\mathbf{w}^{sym}:\mathbb{R}^{D}\to\mathbb{R}^{C}$ be a vector field. Assume that $\mathbf{w}^{sym}$ is rotationally covariant around a point $\mathbf{x}_{0}\in\mathbb{R}^{D}$. Consider an isotropic radially symmetric grid centered at $\mathbf{x}_{0}$, discretized into $N=A^{D}$ points, where $A$ is the number of radial divisions. The grid points are specified by radial indices $a\in\{1,\dots,A\}$ and angular indices $\boldsymbol{\theta}$. Define the quantum state
|$\ket{\psi}=\frac{1}{\sqrt{\mathcal{N}}}\sum_{c=1}^{C}\sum_{a=1}^{A}\sum_{\boldsymbol{\theta}}w^{sym}_{c}\big{(}\mathbf{x}_{a,\boldsymbol{\theta}}\big{)}\ket{c}\otimes\ket{a}\otimes\ket{\boldsymbol{\theta}},$ (35)
@@
where $w^{sym}_{c}\big{(}\mathbf{x}_{a,\boldsymbol{\theta}}\big{)}$ is the $c^{\text{th}}$ component of $\mathbf{w}^{sym}$ at grid point $\mathbf{x}_{a,\boldsymbol{\theta}}$, and $\mathcal{N}=\sum_{c,a,\boldsymbol{\theta}}\left|w^{sym}_{c}\big{(}\mathbf{x}_{a,\boldsymbol{\theta}}\big{)}\right|^{2}$ ensures normalization. This state can be initialized with $\mathcal{O}(N^{1/D})$ classical evaluations of $\mathbf{w}^{sym}$ and $\tilde{\mathcal{O}}(CN^{1/D})$ quantum gates, where the tilde hides polylogarithmic factors.
We consider a spatial source distribution $\pmb{\chi}=\pmb{\chi}^{sym}$ for which an arbitrary rotation in space and the same rotation of the source vector yield the same result. Formally, let there be a $\mathbf{R}(\mathbf{S})\in\mathrm{SO}(C)$, where $\mathbf{S}\in\mathrm{SO}(D)$; then, rotationally covariant sources satisfy $\mathbf{R}(\mathbf{S})\pmb{\chi}^{sym}(\mathbf{x})=\pmb{\chi}^{sym}(\mathbf{S}\mathbf{x})\;\forall\mathbf{x},\mathbf{S}$. We further denote the rotationally covariant Green’s function that solves the wave equation (1) in a homogeneous medium for the source $\hat{\mathbf{s}}_{\delta}(\mathbf{x},t)=\pmb{\chi}^{sym}(\mathbf{x})\delta(t)$ as $\hat{\mathbf{G}}^{sym}$. In this case, the wave field arising from the source time function $f$ is given by $\hat{\mathbf{w}}^{sym}=\hat{\mathbf{G}}^{sym}*f$. It can be discretized as $\hat{\mathbf{w}}^{sym}\rightarrow\mathbf{w}^{sym}$ and efficiently initialized on a spherical grid centered around the source location using Lemma 3.
Figure 1: Algorithm implementing a point source function on a QC for a 1D acoustic problem. $A$ shows the source-time function for a single point location. We decompose the source-time function in time ($B$), which gives a set of functions that, in superposition, equate to the original source-time function ($C$). The inset at the bottom shows the acoustic parameters profiles with the homogenous volume around the source. By convolving them in time with the homogeneous Green’s function ($D$), we obtain multiple sets of initial conditions at different global times ($E$). We evolve these initial conditions over different time intervals to the final simulation time $T$ ($F$). Here, the superposition of the source slices is equal to the solution of the wave equation ($G$). We can utilize this to efficiently estimate an $l_{2}$-distance to a target state using Theorem 2.
Now let us assume that the medium is generally heterogeneous but there exists a small homogeneous volume around the source with radius $r_{s}$. This may hold true in certain experimental setups and can otherwise be taken as an approximation. Let the travel time in the volume be $T^{hom}_{s}=r_{s}/c^{hom}$ where $c^{hom}$ is the wave speed in the homogeneous medium. If the source time function is active for time period $T_{s}>T^{hom}_{s}$ then we can split the source in $J\approx T_{s}/T^{hom}_{s}$ contributions:
$f(t)\approx\sum_{j=1}^{J}W_{j}(t-\tau_{j})f(t),$ (36)
where $W_{j}$ are sigmoid window functions (see Section B.2 for details). All $J$ source terms for $f(t)$ can be recognized as individual sources that satisfy $\Delta T_{j}\leq T_{j}^{hom}$, where $\Delta T_{j}$ is the duration of the $j^{th}$ source. This means that for none of them, the corresponding wave leaves the homogeneous domain before it terminates. Therefore, the resulting $J$ initial wave fields are all rotationally covariant
$\hat{\mathbf{w}}^{sym}_{j}(\mathbf{x},t-\tau_{j})$ $=\hat{\mathbf{G}}^{sym}(\mathbf{x},t)*(W_{j}(t)f(t+\tau_{j})),$ (37)
$|\mathbf{x}|<r,\quad j=1,...,J.$
By using the same discretization as in the homogeneous case, we can initialize the state $[\mathbf{w}^{sym}_{1};...;\mathbf{w}^{sym}_{J}]$ with a runtime complexity of $\tilde{\mathcal{O}}(JCN^{1/D})$. This state can then be synchronized and measured as introduced in Section V.2.2. Hence, if $J=\tilde{\mathcal{O}}(\text{polylog}(N))$ the runtime advantage remains. The whole workflow is depicted in Figure 1.
Note, that the rotationally covariant initial wave field must be pre-computed classically along a reference direction on the spherical grid with a runtime complexity of at worst $\tilde{\mathcal{O}}(CN^{1/D})$. This is, for example, possible via analytic Green’s functions for acoustic and elastic isotropic (explosive) point sources.
VI Limitations
@@
VI.1 Loss functions
We identify two limitations within the loss function estimations introduced:
(i) Estimating time-integrated loss functions: Currently, our algorithm estimates the loss at a single measurement time step. Computing time-integrated loss functions involves separate simulation experiments for each measurement time, with the overall loss metric computed across all steps using either Theorem 2 or classical post-processing. However, this approach becomes inefficient for long comparison/integration times, where the number of comparison time steps scales worse than $\mathcal{O}(\text{polylog}(N))$. Non-destructive estimation of the loss function [58, 59] could provide a solution by preserving the state throughout the measurement operation, enabling iterative evolution and loss estimation of the wave field for multiple time points.
(ii) Globality of measurement quantities: A limitation of QC loss estimation via observable measurements is the requirement for sufficiently global measurement quantities. Suppose a state is measured in a subspace with only a small probability amplitude fraction compared to the entire quantum state. In that case, more samples are needed to estimate the norm accurately, as the expected value becomes vanishingly small. E.g., ensuring the subspace size is large with respect to $N$ can mitigate this issue in practice.
VI.2 Source Terms
@@
(ii) Situations where the number of initialized nonzero values grows polynomially with $N$ must be avoided to prevent restricting the overall runtime. An example of this could be a continuously active source within an inhomogeneous medium.
(iii) For symmetric sources without compact time support, the requirement for radially symmetric grids may be practically limiting as domains can be non-symmetric around the source location.
VI.3 Damping
A significant limitation of the presented framework is its inability to implement damping due to its non-unitary nature. However, damping in the context of HS has been recently studied, demonstrating that quantum speed-up can be preserved under certain conditions [60, 61]. An alternative approach is to map a damped classical wave equation to an open quantum system. Simulating such systems on QC is a recent research focus [62]. We believe that our work can eventually be combined with these approaches.
VI.4 Time-Dependent Media
A final remark concerns time-dependent media, such as those in time-modulated meta-materials [e.g., 63, 64, 65]. In these cases, $\mathbf{B}=\mathbf{B}(t)$ becomes time-dependent, implying that it no longer commutes with the time derivative. Consequently, a time-dependent symmetric problem is mapped to a time-dependent non-Hermitian Schrödinger equation, rendering the introduced framework inapplicable [66].
VII Code
We present an educational Git repository, where we implement all the presented methods and concepts for the acoustic wave equation and Maxwell’s equations. The numerical implementation is based on a staggered grid finite difference method as detailed in Appendix C. The repository can be found here: github.com/malteschade/Quantum-Wave-Simulation-with-Sources-and-Loss-Functions.
VIII Conclusions
@@
Our method is compatible with any numerical approach that respects the anti-Hermitian nature of the operator and results in a diagonal or block diagonal representation of the material properties. The introduced representation of the wave equations naturally leads to a quantum encoding that enables the measurement of energies in sub-volumes of the model with optimal precision scaling in the number of measurements taken. Additionally, we have shown that the $l_{2}$-distance of the final simulated wave field to a target or observed wave field can be measured with optimal precision scaling and that a broad class of source terms can be implemented efficiently.
Recent studies have indicated that HS could be feasible on noisy QC, potentially preceding the advent of noise-free quantum hardware [67, 68]. Furthermore, new advances in quantum error correction hint that noise-free QC might be realized soon [69]. These findings suggest that our proposed framework may have practical applications in the near future, thereby underscoring the importance of continued research in this field.
@@ Acknowledgements.
We would like to thank Marion Dugué for insights on the source decompositions. Furthermore, we are thankful to Patrick Marty and Dr. Ines Ulrich for sharing insights on acoustic wave physics in the context of medical imaging and to Dr. Václav Hapla for initial brainstorming sessions.
Grant Acknowledgment: C.B. was supported by the Swiss National Science Foundation through a Postdoc.Mobility Fellowship. S.K. was partially supported by the project DT-GEO, grant number 101058129. Furthermore, we gratefully acknowledge the support provided by Google Quantum AI, which has significantly contributed to the advancement of this research.
@@
Author contribution statement: study conception: C. Bösch; sources: M. Schade and C. Bösch; loss functions and energy estimation: M. Schade and C. Bösch, S. Keating; acoustic FD implementation: G. Aloisi; general boundary conditions and constraints: G. Aloisi and C. Bösch; manuscript preparation: M. Schade, C. Bösch, G. Aloisi, S. Keating, A. Fichtner. All authors reviewed the results and approved the final version of the manuscript.